3 brothers dispose a sum of money at the bank.The first one disposes 1/2 of the sum, the second 1/5 of the sum and the last one 3/10.After that the first one takes 2/3 of his sum,the second 1/2 of this sum and the last one 5/12 of his sum ,530$ remain in there account. Find :
a)What sum of money each brother disposes at the bank ?
b)What sum of money remains in the bank for each brother?
PLZZZZZZZZ HELPPPPPPPP FOR GOD SAKEEEEEEE THANKS

Answer by checkley71(8403)

.5X+.2X+.3X=X PERCENTAGE OF THE AMOUNT EACH DEPOSITED.
(2/3).5X+(1/2).2X+(5/12).3X=530 AMOUNT LEFT AFTER THE WITHDRAWELS.
1X/3+.1X+1.5X/12=530
(4*1X+12*.1X+1.5X)/12=530
(4X+1.2X+1.5X)/12=530
6.7/12=530
6.7=12*530
6.7X=6360
X=6360/6.7
X=949.25 IS THE TOTAL AMOUNT DEPOSITED.
949.25*.5=474.625 AMOUNT DEPOSITED BY THE FIRST BROTHER.
949.25*.2=189.85 AMOUNT DEPOSITED BY THE SECOND BROTHER.
949.25*.3=284.775 AMOUNT DEPOSITED BY THE THIRD BROTHER.
474.625*2/3=316.42 FIRST BROTHER REMAINING FUNDS.
189.85*1/2=94.92 SECOND BROTHER REMAINING FUNDS.
284.775*5/12=118.66 THIRD BROTHERS REMAINING FUNDS.